What do the lines with the equations:
x+2y=6
2x=-4y+10
y = -1/2x + 5
Answers to Choose From:
They are perpendicular.
They are all horizontal.
They are all vertical.
They each have a y-intercept of 0.
These are all the same line.
They are parallel.

Answer by ewatrrr(24785)

 
Hi,

 Lines:  

 x + 2y=6           0r    y = -x/2 + 3
2x = -4y + 10       0r    y = -x/2 + 10/4
y = (-1/2)x + 5     0r    y =  (-1/2)x + 5 

  Lines are parallel:  might find them


Wish You the Best in your Studies.